OpenStudy (anonymous):

help please solve the eequation. Show your work 5d=17.1

zepdrix (zepdrix):

\[\large 5d=17.1\]Dividing both sides by 5 gives us,\[\large \frac{5d}{5}=\frac{17.1}{5}\]On the left the 5's will "cancel out". Because a number divided by itself is always 1.\[\large \frac{\cancel5d}{\cancel5}=\frac{17.1}{5}\]Which gives us,\[\large d=\frac{17.1}{5}\]Are you allowed to use a calculator for this problem? :)

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

Cool, that would be your last step then. Plugging that into the calculator to solve for d! c:

OpenStudy (anonymous):

Wow you are a life saver thanks

OpenStudy (anonymous):

Hey just to be sure does d=3.42